40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Nov 15 12:25:38 2011, kiwamu, Update, Green Locking, Y arm ALS : beat-note free run fluctuation noise_budget.png
    Reply  Wed Nov 16 17:29:46 2011, kiwamu, Update, Green Locking, Some updates on the Y end green PDH Yarm_fluctuation.png
       Reply  Thu Nov 17 17:21:22 2011, kiwamu, Update, Green Locking, Y end green PDH servo : it's okay servoTF.pngservofilter.pngerr_suppression.png
          Reply  Fri Nov 18 11:28:39 2011, rana, Update, Green Locking, Y end green PDH servo : it's okay 
             Reply  Fri Nov 18 12:11:24 2011, Zach, Update, Green Locking, Y end green PDH servo : it's okay 
    Reply  Thu Nov 17 18:20:26 2011, kiwamu, Update, Green Locking, Y arm ALS : 1st trial of noise budget Yarm_ALS_2011Nov16.png
       Reply  Mon Nov 21 16:08:04 2011, kiwamu, Update, Green Locking, 2nd trial of Y arm ALS noise budget : broad band noise gone Yarm_ALS_2011Nov19.png
          Reply  Tue Nov 22 15:18:18 2011, kiwamu, Update, Green Locking, broad band noise depends on the gain of Y green PDH. and comaprator broken 
Message ID: 5894     Entry time: Tue Nov 15 12:25:38 2011     Reply to this: 5914   5930
Author: kiwamu 
Type: Update 
Category: Green Locking 
Subject: Y arm ALS : beat-note free run fluctuation 

Locking activity last night :

  The free run beat-note in 532 nm has been measured.

However I couldn't close the ALS loop somehow.

Every time I tried closing the loop it broke the Y end PDH lock in a couple of minutes.

 

noise_budget.png

 

 (Things to be done)

   1.  Optimization of the Y end PDH servo loop

      1.1 Measurement of the arm fluctuation => to allow re-designing the servo shape
      1.2 Preparation of PDH box, and temporary SR560 servo
      1.3 Sanity checks on the modulation depth, reflectivity, PD dark noise and etc.,
      1.4 Make the servo more robust
      1.5 Some modifications on the medm screens
      1.6 Activation of the temperature feedback through the realtime digital control

   2. Refinement of the broadband RFPD setup

      2.1 Investigation of the peak source => there was a relatively big peak around 50 MHz or so.
      2.2 Noise characterization of the frequency detection system
      2.3 Nicer routing of some cables.
      2.4 Make two-more ADC channel connectors
      2.5 Power budget on the PSL beat-note setup => estimate the expected RF level of the beat-note
      2.6 Realignment of the PSL doubling and resetting of the doubling oven temperature
     
  3. Noise budgeting
 
     3.1 IR locked condition  => measure the noise in the green beat-note system.
     3.2 ALS engaged condition
          3.2.0 shot noise
          3.2.1 ADC noise
          3.2.2 PD dark noise
          3.2.3 freq. discriminator noise
          3.3.4 DAC noise through the coil-magnet actuators
          3.3.5 End laser suppression
          3.3.6 Intensity noise
          3.3.7 Thermo-elastic noise
          3.3.8 Thermo-refractive noise

 

ELOG V3.1.3-